# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v4

@@ IV-G2 Evaluation on Deepfake Datasets
As we mention before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [1]. We report the results in Table XIII. Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
@@ IV-H Ablation Study
We conduct an ablation study comparing the performance of different frequency transforms, including Fast Fourier Transform (FFT), Wavelet Transform (Wavelet) and Discrete Cosine Transform (DCT). The results in Table XV shows that FFT, Walevet and DCT all show high attack performance, with DWT performing the best. So we use DCT by default.
@@
We conduct an ablation study in Table XVI to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
@@
In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models. The adversarial examples were crafted from ResNet-50 in Synthetic LSUN(ProGAN) dataset. With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper. As shown in Table XVII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate. Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation. Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.

### v5

@@ IV-G2 Evaluation on Deepfake Datasets
As we mentioned before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [shamshad2023evading], which consists of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [ffhq]. We report the results in Table XIII. Although FPBA is not specifically designed for Deepfake detection, it still achieves the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
@@ IV-H Ablation Study
We conduct an ablation study comparing the performance of different frequency transforms, including Fast Fourier Transform (FFT), Wavelet Transform (Wavelet) and Discrete Cosine Transform (DCT). The results in Table XV show that FFT, Walevet and DCT all show high attack performance, with DWT performing the best. So we use DCT by default.
@@
We conduct an ablation study in Table XVI to investigate the impact of computing the gradient in different domains. In comparison with computing the attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
@@
In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models. The adversarial examples were crafted from ResNet-50 in the Synthetic LSUN(ProGAN) dataset. With the exception of $N$, other hyperparameters keep the same with the default settings in the paper. As shown in Table XVII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate. Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation. Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.
@@ V Conclusion
Acknowledgments
This work was supported in part by the NSF China (No. 62302139, 62576020) and Fundamental Research Funds for the Central Universities of China (PA2025IISL0113, JZ2025HGTB0227).
[Uncaptioned image] Yunfeng Diao is a Lecturer in the School of Computer Science and Information Engineering, Hefei University of Technology, China. He received his PhD from Southwest Jiaotong University, China. His current research interests include computer vision and the security of machine learning. He has published over 30 papers in leading venues such as CVPR, ICLR, ICCV, ICML, IEEE TMM and IEEE TCSVT. He has been an editorial board member of IJAACS, the chief organizer of several workshops at IJCAI, also regularly reviewing for top-tier journals and conferences. He has received the Outstanding Contribution Award at the IJCAI Workshop and the Best Paper Award at the IROS RODGE Workshop.
[Uncaptioned image] Naixin Zhai received the B.S. degree from Hefei University of Technology, Hefei, China, in 2024. He is current working toward the M.S. degree with the Department of Electronic Engineering and Information Science. His research interests include deep learning and synthetic image detection.
[Uncaptioned image] Changtao Miao received his B.S. degree in 2019 from AnHui University. He is currently pursuing the Ph.D. degree in Cyber Science and Technology in University of Science and Technology of China. His research interests include face forgery forensics and face manipulation.
[Uncaptioned image] Zitong Yu (Senior Member, IEEE) received the Ph.D. degree in computer science and engineering from the University of Oulu, Finland, in 2022. Currently, he is an Assistant Professor with Great Bay University, China. He was a Post-Doctoral Researcher with the ROSE Laboratory, Nanyang Technological University. He was a Visiting Scholar with TVG, University of Oxford, from July 2021 to November 2021. His research interests include human-centric computer vision and biometric security. He was a recipient of IAPR Best Student Paper Award, IEEE Finland Section Best Student Conference Paper Award 2020.
[Uncaptioned image] Xingxing Wei received his Ph.D degree in computer science from Tianjin University, and B.S. degree in Automation from Beihang University (BUAA), China. He is now an Associate Professor at Beihang University (BUAA). His research interests include computer vision, adversarial machine learning and its applications to multimedia content analysis. He is the author of referred journals and conferences in IEEE TPAMI, TMM, TCYB, TGRS, IJCV, PR, CVIU, CVPR, ICCV, ECCV, ACMMM, AAAI, IJCAI etc.
[Uncaptioned image] Xun Yang (Senior Member, IEEE) received the Ph.D. degree from the Hefei University of Technology, Hefei, China, in 2017. He is currently a Professor with the Department of Electronic Engineering and Information Science, University of Science and Technology of China (USTC). From 2015 to 2017, he visited the University of Technology Sydney (UTS), Australia, as a Joint Ph.D. Student. He was a Research Fellow with the NExT++ Research Center, National University of Singapore (NUS), from 2018 to 2021. His current research interests include information retrieval, cross-media analysis and reasoning, and computer vision. He serves as an Associate Editor for IEEE TRANSACTIONS ON BIG DATA and Multimedia Systems journal.
[Uncaptioned image] Meng Wang (Fellow, IEEE) is a professor at Hefei University of Technology, China. His current research interests include multimedia content analysis, computer vision, and pattern recognition. He received paper prizes or awards from ACM MM 2009 (Best Paper Award), ACM MM 2010 (Best Paper Award), MMM 2010 (Best Paper Award), SIGIR 2015 (Best Paper Honorable Mention), IEEE TMM 2015 and 2016 (Prize Paper Award Honorable Mention) and ACM TOMM 2018 (Nicolas D. Georganas Best Paper Award), etc. He currently serves on the editorial/advisory boards of IEEE TPAMI, IEEE TMM, IEEE TNNLS, etc. He is a fellow of IEEE and IAPR.
Vulnerabilities in AI-generated Image Detection:
The Challenge of Adversarial Attacks
—Supplementary Document—
Figure 10: Visualization of the spectrum saliency map with different models.
VI Additional Experiments
VI-A Experiment Details
We compared a series of distinguished backbone architectures pretrained on the ImageNet dataset. Specifically, we evaluated ResNet-50, DenseNet-121, EfficientNet-B4, MobileNet-V2, ViT-B/16, and Swin-B. Additionally, we examined Spec equipped with ResNet-34 architecture, as well as DCTA equipped with ResNet-50 architecture. In Table XVIII, we enumerate the accuracy of all models on the test set as observed in our experiments. The meticulous evaluation of these foundational models not only revealed the potential vulnerabilities of AIGI detection models but also highlighted the adversarial threats they face. Our analytical results offer significant insights for understanding and improving the robustness of AIGI detection models.
TABLE XVIII: The accuracy of each detector on different datasets.
Model ResNet DenNet EffNet MobNet Spec DCTA ViT Swin
LSUN(ProGAN) 99.99 99.91 99.96 99.93 99.90 99.91 99.74 100.00
GenImage(SD) 99.55 99.92 99.9 99.93 99.12 99.17 99.48 99.97
FFHQ(Style GAN) 99.81 99.89 99.83 99.63 96.51 96.6 99.55 99.99
VI-B Evaluation on ResNeXt
We use ResNext [xie2017aggregated] as the surrogate to evaluate the attack performance. As reported in Table XIX, FPBA still achieves the best average attack success rate, which is consistent with other surrogate models.
TABLE XIX: Attack evaluation using ResNeXt as surrogate model on the ProGAN-LSUN dataset.
Attack ResNeXt ResNet DenNet Spec DCTA ViT Swin Ave.
MI-FGSM 54.2 29.7 21.3 24.8 49.7 14.2 6.6 28.6
PGD 77.1 35.4 30.9 28.0 50.4 20.6 28.0 38.6
S2I 98.6 90.6 86.9 13.4 36.9 15.0 68.0 58.5
FPBA 99.9 89.0 89.2 21.5 49.0 29.6 64.1 63.2
VI-C Study of Attack Strength
As shown in Table XX, we report the comparison of the attack success rates of our FPBA with existing methods under different attack strengths. The results reflect a pattern consistent with when $\epsilon$=8/255, where our method can achieve globally optimal results compared to previous methods, and a larger perturbation budget consistently enhances adversarial transferability across various surrogate models. However, as the attack strength increases, the noise added causes more and more damage to the image, and the visual impact becomes greater, which is revealed in Fig. 11.
VI-D Spectrum Saliency Map
In Fig. 10, we present the spectrum saliency map of all detectors for the original samples and the spectrum saliency map after frequency domain transformation.
TABLE $X$: The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN(ProGAN) with $\epsilon$=4/255, 8/255 or 12/255. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
IFGSM 52.1 44.6 14.0 36.8 15.0 41.2 12.9 24.6 27.0
MIFGSM 52.1 50.0 23.5 45.4 15.0 45.1 22.9 38.3 34.3
ResNet-50 PGD 53.5 46.3 15.5 38.7 20.3 44.4 12.4 26.0 29.1
S2I 97.1 68.1 21.1 33.0 7.0 18.3 2.0 17.9 23.0
$\epsilon$=4/255 FPBA(Ours) 98.9 93.7 55.0 79.8 15.8 40.8 18.6 75.1 54.1
IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
ResNet-50 PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
$\epsilon$=8/255 FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
IFGSM 52.1 51.9 28.7 50.0 18.3 49.1 40.6 50.2 41.2
MIFGSM 52.1 51.8 28.7 47.4 37.2 50.0 34.7 38.9 41.2
ResNet-50 PGD 96.3 96.1 67.0 91.9 30.6 48.8 58.6 87.5 68.6
S2I 97.3 88.4 78.5 89.3 28.5 47.6 25.4 85.3 63.3
$\epsilon$=12/255 FPBA(Ours) 98.9 99.3 82.9 98.5 20.7 49.2 68.7 98.0 73.9
IFGSM 4.0 5.6 12.5 75.4 12.3 24.6 4.8 3.4 9.6
MIFGSM 9.0 11.5 16.5 75.4 13.7 26.4 6.4 10.9 13.5
MobileNet-V2 PGD 3.4 5.0 11.8 87.7 15.1 25.7 4.5 3.2 9.8
S2I 6.0 7.0 19.2 93.6 5.9 8.0 1.2 5.6 7.6
$\epsilon$=4/255 FPBA(Ours) 7.9 11.8 26.0 99.5 14.4 25.8 4.5 9.5 14.3
IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
MobileNet-V2 PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
$\epsilon$=8/255 FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
IFGSM 18.8 21.4 27.7 75.4 18.0 40.2 11.4 23.0 22.9
MIFGSM 25.2 25.2 33.7 75.4 22.5 48.7 21.6 25.2 28.9
MobileNet-V2 PGD 34.8 39.8 27.6 99.8 35.4 48.5 14.7 45.1 35.1
S2I 28.1 33.5 44.0 98.6 23.3 34.5 7.9 35.5 29.5
$\epsilon$=12/255 FPBA(Ours) 42.3 45.7 56.3 99.9 27.0 42.0 18.7 49.0 40.1
Figure 11: The visual results of adversarial examples with different perturbation budgets.